4(x − 1) − (x + 2) = 12
1
Expert's answer
2013-03-27T14:01:22-0400
4(x − 1) − (x + 2) = 12

First, let's open the brackets:

4x − 4 − x - 2 = 12.

Now let's reduce similar terms

3x − 6 = 12

and transfer the constant term (-6) to the right side:

3x = 18.

Dividing the obtained equation by 3 we'll get the root:

x = 6.